OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A.(–5, –2) B.(6, 0) C.(0, 6) D.(–2, –5) @iGreen

OpenStudy (anonymous):

Hi! This is a plug and chug equation meaning you can just plug in the answers to the equation. Slow but reliable. SO for A it would be 4(-5)-5(-2) can you solve this and tell me your answer?

OpenStudy (igreen):

4a – 5b = 24 Plug in (-5, -2) into the equation: 4(-5) - 5(-2) = 24 Multiply: -20 + 10 = 24 Add: -10 = 24 (-5, -2) is not a solution to the equation. Plug in (6, 0): 4a - 5b = 24 4(6) - 5(0) = 24 Multiply: 24 - 0 = 24 Can you subtract 24 - 0? @NunChuckBoy3101

OpenStudy (anonymous):

Never mind. Igreen knows how to solve this far better then me :D

OpenStudy (anonymous):

24-0=24

OpenStudy (anonymous):

ok thank you for trying @ohlookidontunderstandanotherthing

OpenStudy (igreen):

Right, so that leaves you with 24 = 24, which means (6, 0) is a solution to 4a - 5b = 24.

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

@iGreen thank you if i have any more questions can i tag you?

OpenStudy (igreen):

Yes, you can.

OpenStudy (igreen):

And you're correct, B is the right answer.

OpenStudy (anonymous):

thank you